In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?
The answer is 17, but how do you get this answer? I don't know what formula to use to solve this

Answer by Nihal@SriLanka(22)

Your notion that there is a formula for the solution of such problems is wrong.

What we do is we express the given information as a simple algebraic equation with one unknown figure ( In this case the no of pages of smaller type. Note that once you know the pages of smaller type then the pages with large type is 21 less that no )

We denote no of small type pages by x representing a no and the equation turns out to be

x * 2400 + (21-x) * 1800 = 48000 which we obtain by simply adding the words in all 21 pages assuming small type page no to be x.
progressively simplifying above equation we have
2400x + 21 * 1800 - 1800x = 48000
i.e. 600x = 48000 - 21 * 1800 = 48000 - 37800 = 10200
i.e. x = 17
If you need any further clarification you may email sumanapala@gmail.com